You are standing still. Light explodes at 600m away and reaches us 3micro seconds, 2 micro seconds later another light explosion from 1200m away reaches you. Which firecracker explodes first and at what micro seconds did they explode?
1
Expert's answer
2017-02-06T08:38:11-0500
L1= 600m; t1=3*10^-6 s
L2=1200m; t2= 5*10^-6 s;
v1=v2=c= 3*10^8 m/s; (speed of light)
T1= L1/v1= 600/3*10^8= 2*10^-6 s
T2=L2/v2=1200/3*10^8=4*10^-6 s
Time of explosions: t10= t1-T1= 10^-6s= 1 microsecond (explodes first firecracker)
t20= t2-T2= 10^-6s= 2 microsecond (explodes second firecracker)
Answer:
Firecracker #1 explodes first at 1microsecond
Firecracker #2 explodes at 2 microsecond